In a clinical t a 27 out of 950 patients taking a presc ption ng compla d on ke symptoms. Suppose that it s known that 1.8% of patients taking c r peting drugs complain of fulike symptoms. Iis there sufficient evidence to conclude that more than 1.8% of this drugs users experience flulike symptoms as a side effect at the a 0.1 level of significance? Choose the correct answer below. OA No, tereis not suficient evidence that more than 1.8% exper encefl kesymptoms, O B. No, there is not suffient evidence that more than 1.8% experience fulike symptorms, because the test statistic is greater than the critical value, meaning that O c Yes, there is sufficient evidence that more than 1.8% experience fulike symptoms, because the test statistic s less than the critical value. meaning that we 0 D Yes, there is suficient evidence that more than 1.8% experience fulke symptoms, because the test statistic s greater than the critical value, meaning that we because the test statistic is less than the critical value, meaning that we do not reject the null hypothesis. we reject the null hypothesis. do not reject the null hypothesis. reject the null hypothesis
Answer #1

n=950

x = 27

Sample proportion, ar p = 27/950 = 0.028

(1) Null and Alternative Hypotheses:

Ho: p = 0.018

Ha: p > 0.018

(2) Critical value:

At alpha = 0.1 right-tailed critical value, zc = 1.28

(3) Test Statistics:

0.028 - 0.018 : = = 2.32 V p(1 - p)/n /0.018(1 0.018)/950

(4) Decision:

As z = 2.32 > zc = 1.28, it is then concluded that the null hypothesis is rejected.

Answer D

Yes, there is sufficient evidence that more than 1.8% experience flulike symptoms, because the test statistic is greater than the critical value, meaning that we reject the null hypothesis.